OpenStudy (hockeychick23):

Mike bought a new computer game at the store. He went home and played the game all day. The list below shows the scores for just the first 10 games that Mike played. 15, 13, 8, 12, 15, 17, 14, 9, 11, 16 If Mike scored 30 on his next game, would it be an outlier for the set of his first 11 games? Explain. How would a score of 30 affect the range, mean, median, and mode of his scores? (I would say yes it is an outlier because 30 is so far away from the median(13.5) and mean(13), and that it would bring the median, range, and mean up but wouldn't effect the mode. Am I right? thanks!

OpenStudy (zzr0ck3r):

I thought there was a definition for outlier.

OpenStudy (hockeychick23):

Yea there is its: a data value that is markedly different from most of the other data values in a sample. I just wanted to make sure my explanation worked and was ok

OpenStudy (ganpat):

Mean of 10 scores = 130/10 = 13 Mean of 11 scores = 160/11 = 14.54 your attempt was right, but the difference is not that extreme.

OpenStudy (hockeychick23):

am i looking at the difference in the means though or the difference between 31 and the other numbers in the sample?

OpenStudy (hockeychick23):

*30

OpenStudy (ganpat):

just 30 as score for the 11th game seems to be an outlier But its effect on other factors need to be calculated as well !! How would a score of 30 affect the range, mean, median, and mode of his scores? soo you have to calculate range, mean and median..

OpenStudy (ganpat):

range shall be affected as well.. but not mean, median and mode.

OpenStudy (hockeychick23):

yea i said the range would have a big change because now instead of being a range of 9 it would be a range of 22

OpenStudy (ganpat):

true that..

OpenStudy (aum):

If you want a mathematical way to determine if a data point is an outlier you can use this method: http://www.wikihow.com/Calculate-Outliers